The denominator of a fraction is two more than the numerator. If both numerator and denominator are increased by ten, the simplified result is 9/10. Find the original fraction. (Do NOT simplify.)

Answer by Boreal(15235)

let the fraction be (x)/(x+2)
the new one is (x+10)/(x+12)=9/10
cross-multiply
9(x+12)=10(x+10)
9x+108=10x+100
x=8
x+2=10
(8/10) is the fraction